OpenStudy (anonymous):

What is the area of a circle with diameter 12 inches? 144 in 2 6 in 2 12 in 2 36 in 2

OpenStudy (anonymous):

36 pi

OpenStudy (anonymous):

a i guess

OpenStudy (akshay_budhkar):

36pi

OpenStudy (anonymous):

area = pi r^2 r = half diameter

OpenStudy (anonymous):

None 36 Pi

OpenStudy (anonymous):

r=d/2 r=12/2 r=6 area=pi*r^2